A roof is shaped like a rectangle with a length that is twice the width and an area of 1800 sq ft. What is the lenght, width and perimeter of the roof?
Answer by rajagopalan(174)

A roof is shaped like a rectangle.
Let the width=x
Length=2x
Area=2x^2=1800 sq ft.
2x^2=1800
x^2=1800/2=900
x=sqrt900=30
Width=30
Length=2*30=60
perimeter=2(x+2x)
=2(30+60)
=2(90)
=180 ft